A 0.5kg basketball moving at 1.4m/s collides with a 1.2kg bowling ball moving at 3.5m/s. After collision, the velocity of the basketball was increased by 0.4m/s. What is the final velocity of the bowling ball?

1
Expert's answer
2022-03-16T09:41:58-0400

Solution.

"m_1=0.5 kg;"

"v_1=1.4 m\/s;"

"m_2=1.2 kg;"

"v_2=3.5 m\/s;"

"u_1=1.8 m\/s;"

"u_2 -?;"

"m_1v_1-m_2v_2=-m_2u_2-m_1u_1;"

"m_2u_2=-m_1v_1+m_2v_2-m_1u_1;"

"u_2=\dfrac{-m_1v_1+m_2v_2-m_1u_1}{m_2};"

"u_2=\dfrac{-0.5\sdot1.4+1.2\sdot3.5-0.5\sdot1.8}{1.2}=2.2 m/s;"

Answer: "u_2=2.2 m\/s."